OpenStudy (chris215):

Which of the following is the best linear approximation for f(x) = sin(x) near x = π seconds? y = -x + π - 1 y = -1 y = -x + π y = -x - π

zepdrix (zepdrix):

Hmm I don't think you need to do anything fancy here. Just think about the `actual value` of sin(x) at x=pi, and see which options gets you closest to that value.

zepdrix (zepdrix):

\[\large\rm \sin(\pi)=?\]

OpenStudy (anonymous):

@zepdrix is right. TIP: It's equally beneficial to glance at a sine wave on a graph.

OpenStudy (anonymous):

For a more systematic way (assuming you know calculus), the linear approximation of a function at a point is: \(L = f(x_{0}) + f'(x_{0})(x-x_{0})\) Then just throw everything into that formula. \(\large \sin(\pi) = ?\) \(\large \frac{d}{dx}\sin(x) \rvert_{x=\pi} = ?\)

OpenStudy (chris215):

-1?

zepdrix (zepdrix):

\(\large\rm \sin(\pi)\ne-1\)

OpenStudy (chris215):

ok so does that mean the answer is y=-1?